What are the measures of a linear pair of angles if the measure of one angle is five times that of the other?

Answer by stanbon(75887)

What are the measures of a linear pair of angles if the measure of one angle is five times that of the other?
---
Equations:
x + y = 180
x = 5y
-----
Substitute for "x" and solve for "y":
---
5y + y = 180
6y = 180
y = 30 degrees
----
Solve for "x":
x = 5y
x = 5*30 = 150 degrees
================================
Cheers,
Stan H.